# Finite groups and arc-transitive maps of square-free Euler characteristic

### v2

@@ 4. $p$-Groups
which gives an expected factorization of $G$. We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm
Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $A_{r}$ of $A$, so that, $G_{p}$ is centralized by $A$.
By Lemma 4.3, we have two cases to treat: $A_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
Case 1. Let $A_{r}={\sf Z}_{r^{k}}$. If $A_{r},B_{r^{\prime}}]=1$, then by Lemma 4.3, $(A{:}B)_{r}=A_{r}{:}B_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian, and
$1\not=A_{r}\cap{\bf Z}(A_{r}{:}B_{r})=A_{r}\cap{\bf Z}(A{:}B)\leqslant A\cap{\bf Z}(A{:}B),$
@@
a contradiction. Thus there exists a Sylow $q$-subgroup $B_{q}$ with $q\not=r$ which does not centralize $A_{r}$. Take $x\in B_{q}$ to be an element acting non-trivially on $A_{r}$ and let $X=A_{r}{:}\langle x\rangle$. Then $X$ acts on $G_{p}$ and induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. It then follows by Lemma 5.3 (i) that $A_{r}\leqslant X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
@@
Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (2), it implies that either
  • (i)
    $A_{r}={\sf Z}_{r}^{2}$ and $B$ is irreducible on $A_{r}$, or
  • (ii)
@@
    there exists $a,b\in A_{r}$ such that $A_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
@@
For the former, $A_{r}$ centralizes $G_{p}$ since $A_{r}{:}B$ induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. For the latter, there exist elements $u,v\in B$ such that $u$ acts on $\langle a\rangle$ non-trivially, and $v$ acts on $\langle b\rangle$ non-trivially. Then similar to Case 1 we have $G_{p},a]=[G_{p},b]=1$, so that, $G_{p},A_{r}]=[G_{p},\langle a,b\rangle]=1$.
We have therefore confirmed the claim. Consequently we have
$G=G_{p}{:}(A{:}B)=(G_{p}\times A){:}B.$
Meanwhile, by Lemma 4.3, $G_{p}$ is one of the following groups:
$G_{p}={\sf Z}_{p^{\ell}}$,  ${\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$  or  ${\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$.
Note that by assumption, $G_{p^{\prime}}=A{:}B$ acts by conjugation on $G_{p}$ non-trivially, which implies that $B$ acts on $G_{p}$ non-trivially.
@@
Suppose $G_{p}={\sf Z}_{p^{\ell}}$. As $B$ acts on $G_{p}$ non-trivially, it follows from Lemma 5.3 (ii) that $G_{p}\cap{\bf Z}(G)=1$. Thus G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$. By Lemma 4.3 (3), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially. Rewrite $G$ as
$G=(A\times(\langle a\rangle{:}\langle b\rangle)){:}B=(A\times\langle a\rangle){:}(B\times\langle b\rangle).$
Then by Lemma 5.3 (ii), G=C:DG=C{:}D satisfies the theorem, with $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$.
@@
Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$. By Lemma 4.3 (2), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$. For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$. For the latter, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$. Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$. Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$. $\Box$
@@
Lemma 5.5.
Let $G$ be a solvable group satisfying Hypothesis 2.2. If $F:={\rm Fit}(G)={\bf O}_{2}(G)$, then either $G$ is a $2$-group, or $G$ is one of the following:
  • (1)
    $G={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ or ${\sf Z}_{2}^{3}{:}{\sf Z}_{7}{:}{\sf Z}_{3};$
  • (2)
    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
Proof.
Assume that $G$ is not a 2-group. Then $G/F\leqslant{\sf Out}(F)$ is not a 2-group either. Thus by Lemma 4.1, $F$ is one of the following 2-groups:
$F\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}.$
Moreover, ${\bf O}_{2}(G/F)$ is trivial, namely, $G/F$ has no normal 2-subgroups. Then by checking ${\sf Out}(F)$, either $|G/F|$ is divisible by 7, and $F={\sf Z}_{2}^{3}$, $G/F={\sf Z}_{7}$ or ${\sf Z}_{7}{:}{\sf Z}_{3}$; or $G/F={\sf Z}_{3}$ or ${\rm S}_{3}$. It follows that $G$ is one of the groups listed in the lemma. In particular, if $G=F{:}{\rm S}_{3}$, then $G_{2}=F{:}{\sf Z}_{2}>F$, and respectively for $F={\sf Z}_{2}^{2}$, ${\sf Z}_{2}^{3}$, ${\rm Q}_{8}$ or ${\sf Z}_{4}\circ{\rm Q}_{8}$, it is easy to verify that $G_{2}$ has a subgroup $K$ of index $2$ with $K={\sf Z}_{4}$, ${\rm D}_{8}$, ${\sf Z}_{8}$ or ${\rm D}_{16}$. That is, $G=F{:}{\rm S}_{3}$ surely satisfies Hypothesis 2.2. $\Box$
Lemma 5.6.
Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. If $K\neq K_{p}{:}K_{2}$, then one of the following holds.
@@
    $p=3$, $K=X{:}K_{3}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$
  3. (3)
    $p=3$, $K$ is one of the groups in Tables 1 - 2
Proof.
Suppose that $K_{2}$ is normal in $K$, that is, $K=K_{2}{:}K_{p}$. Then $K_{2}$ has an automorphism of order $p$. By Lemma 4.1, $K_{2}$ is one of the following groups:
${\sf Z}_{2}^{2},\ {\rm Q}_{8},\ {\sf Z}_{2}^{3}\ \mbox{and}\ {\sf Z}_{4}\circ{\rm Q}_{8},$
and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in (1) - (2).
We then assume that neither $K_{p}$ nor $K_{2}$ is normal in $K$. Let
$F={\rm Fit}(K)=F_{p}\times F_{2}$, with $F_{p}<K_{p}$, $F_{2}<K_{2}$.
Consider the quotient group $\overline{K}=K/F_{p}$, whose Fitting group is a 2-group, namely ${\rm Fit}(\overline{K})={\bf O}_{2}(\overline{K})$. Thus, by Lemma 5.5, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$, where either
(i) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (ii) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
Further, if (ii) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity. Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
Therefore, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is given in (i). We then analyze all the possibilities for $F_{2}=\overline{F_{2}}$, which a normal 2-subgroup of $\overline{K}$.
Case 1. $F_{2}={\bf O}_{2}(\overline{K})$. In this case, we have $H={\rm S}_{3}$ since $F_{2}<K_{2}$. Then
$K=F_{3}.(F_{2}{:}{\rm S}_{3})=(F_{3}\times F_{2}).{\rm S}_{3}=F_{2}{:}(F_{3}.{\rm S}_{3})=F_{2}{:}(K_{3}{:}{\sf Z}_{2}).$
Note that $K_{3}{:}{\sf Z}_{2}/F_{3}\cong{\rm S}_{3}$. By checking the $3$-groups appearing in Lemma 4.3, we conclude that there are seven possibilities for the triple $T:=(K_{3},\,K_{3}{:}{\sf Z}_{2},\,F_{3})$ as follows (also note that ${\sf Aut}({\sf Z}_{3}\times{\sf Z}_{3})={\rm GL}_{2}(3)$ is solvable):
[…]
Consequently, Case 1 gives the groups listed in Table 1.
Table 1. Case 1: $F_{2}={\bf O}_{2}(\overline{K})$
case $F_{2}$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
(1.1) ${\rm S}_{4}$ ${\sf Z}_{2}\times{\rm S}_{4}$ ${\rm GL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$
(1.2) ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}})$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}})$
(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
(1.5) $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{2.3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$
(1.6) $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{12}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$
(1.7) ${\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{2}\times{\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}))$
Case 2. $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ or $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$. Then
$K=F_{3}.({\bf O}_{2}(\overline{K}){:}H)=(F_{3}\times F_{2}).({\sf Z}_{2}\times H)=F_{2}{:}(F_{3}.({\sf Z}_{2}\times H)).$
Further, if $H={\sf Z}_{3}$, then $K/F_{2}=K_{3}{:}{\sf Z}_{2}$. Note that ${\sf Z}_{2}$ acts non-trivially on $K_{3}$, and meanwhile it acts trivially on $K_{3}/F_{3}=H$. Then by checking the $3$-groups appearing in Lemma 4.3, we conclude that there are three possibilities:
[…]
Otherwise, $H={\rm S}_{3}$, and so $K/F_{2}=K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})$. By Lemma 4.3, ${\sf Z}_{2}\times{\sf Z}_{2}\leqslant{\sf Aut}(K_{3})$ implies $K_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and $K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$. We then get two possibilities:
[…]
Consequently, Case 2 gives the groups listed in Table 2.
Table 2. Case 2: $F_{2}<{\bf O}_{2}(\overline{K})$
case $(F_{2},{\bf O}_{2}(\overline{K}))$ $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$
(2.1) ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
(2.2) $({\sf Z}_{2}^{2}{:}{{\sf Z}_{3^{\ell}}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{{\sf Z}_{3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
(2.3) $({\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}^{2}){:}{\sf Z}_{3}$ $(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ{\rm Q}_{8}){:}{\sf Z}_{3}$
(2.4) ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(2.5) $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
Case 3. Recall that $F_{2}$ is a normal 2-subgroup of ${\bf O}_{2}(\overline{K})$. Notice that ${\rm Q}_{8}{:}{\sf Z}_{3}$ or ${\rm Q}_{8}{:}{\rm S}_{3}$ only has the center ${\sf Z}_{2}$ and ${\rm Q}_{8}$ as normal $2$-subgroups. Then all the remaining cases are collected as follows :
[…]
For case (3.1), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\rm Q}_{8})$, we have
$F=F_{3}\mbox{\ and\ }K/F_{3}={\rm Q}_{8}{:}H\lesssim{\sf Aut}(F_{3}).$
By lemma 4.3, it implies that $F_{3}={\sf Z}_{3}\times{\sf Z}_{3}$ and $K_{3}={\sf Z}_{9}{:}{\sf Z}_{3}$ is non-abelian. Let
$K_{3}=\langle a\rangle{:}\langle b\rangle={\sf Z}_{9}{:}{\sf Z}_{3}$, with $a^{b}=a^{4}$, $b^{a}=a^{6}b$,
where $F_{3}=\langle a^{3}\rangle\times\langle b\rangle$. Note that $K_{2}=({\rm Q}_{8}{:}H)_{2}={\rm Q}_{8}$ or ${\rm Q}_{8}{:}{\sf Z}_{2}$. Consider the central involution $c$ of $K_{2}$. Then $\overline{c}$ is the central involution of $\overline{K}$. On the one hand, $\overline{c}$ acts on $F_{3}$ by mapping each $x$ to $x^{-1}$. That is,
$x^{c}=x^{-1}$, so that, $c^{x}=x^{-2}c$ for any $x\in F_{3}$.
On the other hand, $\overline{c}^{\overline{a}}=\overline{c}$, and so $c^{a}=cy$ where $y\in F_{3}$. Let $y=a^{3i}b^{j}$. Then
$c^{a^{3}}=(ca^{3i}b^{j})^{a^{2}}=(ca^{(6i+6j)}b^{2j})^{a}=ca^{(9i+18j)}b^{3j}=c,\mbox{\ where\ }a^{3}\in F_{3},$
a contradiction. Hence case (3.1) cannot occur.
For case (3.2), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\sf Z}_{4}\circ{\rm Q}_{8})$, we have ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3})\lesssim{\sf Aut}(F_{3})$, which is impossible by checking the 3-groups appearing in Lemma 4.3.
For case (3.3), since ${\bf O}_{2}(\overline{K})$ acts on $F_{3}$ with the centralizer ${\bf C}_{{\bf O}_{2}(\overline{K})}(F_{3})=F_{2}$, we deduce that in this case ${\sf Aut}(F_{3})$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$. By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by $4$ implies $F_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$. For $\ell=1$, it is clear that ${\sf Aut}({\sf Z}_{3}^{2})={\rm GL}(3,2)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ contains no subgroups isomorphic to ${\rm A}_{4}$. We thus assume $\ell>1$ and let $F_{3}=\langle a\rangle\times\langle b\rangle$ with $|a|=3^{\ell}$, $|b|=3$. By Lemma 4.3, ${\sf Aut}(F_{3})$ is solvable, and it contains a Hall $3^{\prime}$-subgroup
$H_{0}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}\cong{\sf Z}_{2}^{2},$
which stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$. Meanwhile, by simple calculation, $H_{0}$ cannot stabilize
Consider the normalizer $N={\bf N}_{{\sf Aut}(F_{3})}(H_{0})$. For any $g\in N$, $H_{0}=H_{0}^{g}$ stabilizes subgroups $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, and so $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, which implies that $N\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian. Note that any subgroup $H\leqslant{\sf Aut}(F_{3})$ with $H\cong{\sf Z}_{2}^{2}$ is a Hall $3^{\prime}$-subgroup, then conjugate to $H_{0}$. Then the normalizer ${\bf N}_{{\sf Aut}(F_{3})}(H)$, which is conjugate to $N$, is also abelian. Thus, ${\sf Aut}(F_{3})$ has no subgroup isomorphic to ${\rm A}_{4}$.
Finally, with Case 3 excluded, the proof is complete. $\Box$
@@
By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. Thus $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Lemma 5.6,
$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7},$
Let $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$. As $K_{2}$ is normalized by $K_{7}$, and meanwhile it normalizes $F_{7}$, so $K_{2}$ centralizes $F_{7}$. Moreover, $K_{2}$ also centralizes $F_{3}$. If not, consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$. Then ${\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$, which is impossible due to Lemma 4.3 and the fact ${\sf Z}_{2}^{3}\nleqslant{\rm GL}(2,3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$. Hence $K_{2}$ centralizes $F$, so $K_{2}\leqslant C_{K}(F)\leqslant F$, and then $K_{2}=F_{2}\lhd K$, and $K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})$. $\Box$
Now we are ready to prove Theorem 1.1.
@@
Let $G$ be a finite solvable group of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index. Let $H$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$. Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$. Then we have factorization

### v3

@@ 4. $p$-Groups
We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm the following claim.
Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $C_{r}$ of $C$, so that, $G_{p}$ is centralized by $C$.
By Lemma 4.3, we have two cases to treat: $C_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
Case 1. Let $C_{r}={\sf Z}_{r^{k}}$. If $C_{r},D_{r^{\prime}}]=1$, since $C_{r},D]\neq 1$, then $(C{:}D)_{r}=C_{r}{:}D_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian as given in Lemma 4.3 (3), and so
$1\not=C_{r}\cap{\bf Z}(C_{r}{:}D_{r})=C_{r}\cap{\bf Z}(C{:}D)\leqslant C\cap{\bf Z}(C{:}D),$
@@
a contradiction. Thus, there exists a Sylow $q$-subgroup $D_{q}$ with $q\not=r$ which does not centralize $C_{r}$. Let $x\in D_{q}$ act non-trivially on $C_{r}$, and let
$X=C_{r}{:}\langle x\rangle.$
Then, $X$ acts on $G_{p}$ with $\overline{X}=X/{\bf C}_{X}(G_{p})\lesssim{\sf Aut}(G_{p})$. Note that, $|X|$ is coprime to $2p$, so is $|\overline{X}|$. By Lemma 4.3, either ${\sf Aut}(G_{p})={\rm GL}_{2}(p)$, or ${\sf Aut}(G_{p})$ is solvable with an abelian Hall $p^{\prime}$-subgroup, while both cases imply that $\overline{X}$ is abelian, that is, $X^{\prime}\leqslant{\bf C}_{X}(G_{p})$. At last, by Lemma 5.2 (i), we have $C_{r}\leqslant X^{\prime}$, and so $C_{r}\leqslant{\bf C}_{X}(G_{p})$.
@@
Case 2. Let $C_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. In this case, $C_{r}=(C{:}D)_{r}$, namely, $r$ does not divide $|D|$. It follows by Lemma 4.3 (2) that, either
    $C_{r}={\sf Z}_{r}^{2}$ and $D$ is irreducible on $C_{r}$; or
@@
    there exists $a,b\in C_{r}$ such that $C_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $D$.
@@
For the former, let $x\in D$ act non-trivially on ${\sf Z}_{r}^{2}$, and let $X={\sf Z}_{r}^{2}{:}\langle x\rangle$, which has ${\sf Z}_{r}^{2}$ as a minimal normal subgroup; for the latter, there exist $u,v\in D$ respectively acts on $\langle a\rangle$, $\langle b\rangle$ non-trivially, so let $Y=\langle a\rangle{:}\langle u\rangle$, $Z=\langle b\rangle{:}\langle v\rangle$. Note that, each of $X,Y,Z$ induces some automorphism group of $G_{p}$ of order coprime to $2p$. It is similar to Case 1 to show $G_{p},C_{r}]=1$.
We have therefore confirmed the claim. Consequently, we have
$G=G_{p}{:}(C{:}D)=(G_{p}\times C){:}D.$
We then find an expected factorization of $G$. By assumption, $G_{p^{\prime}}=C{:}D$ acts by conjugation on $G_{p}$ non-trivially. Now, as $C$ acts on $G_{p}$ trivially, $D$ acts on $G_{p}$ non-trivially, so it induces some automorphism of $G_{p}$ of order coprime to $p$.
By Lemma 4.3, we have three cases to treat:
@@
(1) Suppose $G_{p}={\sf Z}_{p^{\ell}}$. By Lemma 5.2 (ii), any non-identity element of $G_{p}$ cannot fixed by $D$. Thus G=A:BG=A{:}B, with $A=G_{p}\times C$, $B=D$, satisfies the theorem.
(2) Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ as given in Lemma 4.3 (3). Then there exist $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle{:}\langle b\rangle$, $\langle a\rangle$ is normalized by $D$, and $\langle b\rangle$ is centralized by $D$. Rewrite $G$ as
$G=(C\times(\langle a\rangle{:}\langle b\rangle)){:}D=(C\times\langle a\rangle){:}(D\times\langle b\rangle).$
By Lemma 5.2 (ii), G=A:BG=A{:}B, with $A=C\times\langle a\rangle$, $B=D\times\langle b\rangle$, satisfies the theorem.
@@
(3) Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$. It follows by Lemma 4.3 (2) that, either $G_{p}={\sf Z}_{p}^{2}$ and $D$ is irreducible on $G_{p}$; or there exist $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$, $\langle b\rangle$ are normalized by $D$. For the former, G=A:BG=A{:}B satisfies the theorem, with $A=G_{p}\times C$, $B=D$. For the latter, if $D$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $A=G_{p}\times C$, $B=D$; otherwise, $D$ acts non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$, so let $A=C\times\langle a\rangle$, $B=D\times\langle b\rangle$. $\Box$
@@
Proposition 5.4.
Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. Assume that $K\neq K_{p}{:}K_{2}$. Then $p\in\{3,7\}$ and one of the following holds.
@@
    $K$ is one of the $\{2,3\}$-groups: $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$ and $W$, where $X,Y,W$ are given in Table 1, $X$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\rm S}_{3}\cong{\rm S}_{4}$, $Y$ is homomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}_{2}(3)$, and $W$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\sf Z}_{3}\cong{\rm A}_{4}$, ${\rm S}_{4}$, ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$.
We remark that, the proof of Proposition 5.4 has been moved to the end of this Section, as it is rather lengthy.
@@
By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. If $K_{\{2,7\}}=K_{7}{:}K_{2}$, then $K_{7}$ is normal in $K$, not in our case. Thus, $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Proposition 5.4, we have
$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7}.$
We claim that, $K_{2}={\sf Z}_{2}^{3}$ centralizes the Fitting group $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$. If this is true, then we are done, since $K_{2}\leqslant C_{K}(F)\leqslant F$, so $K_{2}=F_{2}\lhd K$, and
$K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})={\sf Z}_{2}^{3}{:}(K_{7}{:}K_{3}).$
It follows by $F_{2}\leqslant K_{2}={\sf Z}_{2}^{3}$ that $K_{2}$ centralizes $F_{2}$. Meanwhile, $K_{2}$ is normalized by $K_{7}$, so $K_{2}$ and $F_{7}$ normalizes, and hence centralizes, each other. The claim is then proved if $K_{2}$ centralizes $F_{3}$. Suppose that $K_{2}$ doe not centralize $F_{3}$. Consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$. Then $K_{2}={\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$. By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by 8 only happens when $F_{3}={\sf Z}_{3}^{2}$, ${\sf Aut}(F_{3})={\rm GL}_{2}(3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$. However, ${\rm GL}_{2}(3)$ has no subgroup isomorphic to ${\sf Z}_{2}^{3}$, a contradiction. $\Box$
Now, we are ready to prove Theorem 1.1.
@@
Let $G$ be a finite solvable group of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index. Then, the Sylow subgroups $G_{2}$ and $G_{p}$ ($p>2$) are respectively given in Proposition 4.1 and Lemma 4.3.
Let $H\leqslant G$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$. Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$. Then, we have :